OpenStudy (anonymous):

How would I solve x+2y=20, and x-y=16

OpenStudy (jdoe0001):

well, is a system of equations take a peek at the 2nd equation x-y = 16 can you solve that for "x", what would it look like?

OpenStudy (anonymous):

Wouldn't you start out with x-y=16, then convert it to x=-16+y?

OpenStudy (jdoe0001):

well, yes, but what you have there is 16, not -16, so it'd be x-y = 16 x = 16 +y

OpenStudy (jdoe0001):

so we have x+2y=20 x-y=16 and from solving for "x" on the 2nd equation, we now know that x = 16 +y now if I were to rewrite the 1st equation, with x = 16+y in mind, it'd be \(\color{blue}{x}\) + 2y = 20 OR \(\color{blue}{(16+y)}\) + 2y = 20 16 + y + 2y = 20 so now just solve for "y", what would that give you for "y"?